OpenStudy (anonymous):

What is the reverse Laplace Transform of X(s)= (s^2 +6)/[(s)(s+1)(s+3)]?

OpenStudy (anonymous):

Partial Fractions it first. Break the denominator up into three parts

OpenStudy (anonymous):

\[\frac{s^2 + 6}{(s)(s+1)(s+3)} = \frac{A}{s} + \frac{B}{s+1} \frac{C}{s+3}\]

OpenStudy (anonymous):

+ between** B/s+1 + C/s+3 Solve for A B and C

OpenStudy (anonymous):

Thank you, I think i have it now.

OpenStudy (anonymous):

X(s)= 2/s - 2/7(s+1) +2/5(s+3)

OpenStudy (anonymous):

I think it shoud be 7/2 and 5/2

OpenStudy (anonymous):

X(s)= 2/s - 2/[7(s+1)] +2/[5(s+3)]*

OpenStudy (anonymous):

sorry for the late response the attached should be the answer. But I am sure you already got it =]

OpenStudy (anonymous):

Thanx :)